OpenStudy (anonymous):

Ages. Tyler is twice as old as his son. Ten years ago, Tyler was three times as old as his son. How old are they now?

OpenStudy (lalaly):

let stand for Tyler's age and s stand for his son's age. t = 2s --->> It says that Tyler is twice as old as his son. t - 10 = 3(s -10) --- This is because if you take t, Tyler's current age, and subtract 10 from it, you get Tyler's age 10 years ago. However, you must also take those 10 years away from his son, because they would have both been 10 years younger. After that, you just need to multiple his son's age by 3 and set the two expressions equal to each other.

OpenStudy (lalaly):

let t*

OpenStudy (anonymous):

im still confused uhh

OpenStudy (anonymous):

Your variables are t and s. t representing Tyler's age and s representing his son's age. " Tyler is twice as old as his son" ---- This can be translated into t=2s because this literally means Tyler's age is 2 times his son's age. The next line is a little more complicated. "Ten years ago, Tyler was three times as old as his son." Ten years ago Tyler and his son were both ten years younger. So when translating this sentence we'll use (T-10) to represent Tyler's age ten years ago and (s-10) to represent his son's age ten years ago (These expressions literally mean Tyler's age decreased by 10 and his son's age decreased by 10). We know that ten years ago Tyler was three times older than his son. So... Tyler's age ten years ago is 3 times his son's age ten years ago. This can be translated into t-10=3(s-10). We're asked to find their ages. Our two equations are---- t=2s (Tyler's age is twice his son's age) AND t-10=3(s-10) (Ten years ago Tyler was three times older than his son) We need to solve for our variables to find Tyler and his son's ages. We can use substitution to get the equation--- (2s)-10=3(s-10) Notice that since t equals 2s anywhere I see "t" in an equation I can replace it with "2s" since they are equal. Now I can solve this new equation for s (2s)-10=3(s-10) using distributive prop. I get--- 2s-10=3s-30 I'll add 30 to both sides 2s+20=3s Now I'll subtract 2s from both sides isolating my variables I want to solve for (s) s=20 I can substitute 20 into the original equation T=2s to find that T=40. Notice that these ages work. Tyler(40 years old) is twice as old as his son (20 years old) and that ten years ago Tyler would have been thirty while his son would have been ten.

OpenStudy (anonymous):

ty very muchhh

OpenStudy (anonymous):

not a problem let me know if any of it is still confusing.